Fermi National Laboratory
 

Accelerator Update for May 19, 2006 - May 22, 2006

 

Friday May 19

The day shift began with the Booster in access.

Technicians replaced a Booster RF station (BRF6).

Operators and technicians accessed Linac.

The cryo system techs received permission to cooldown the Tevatron (TeV) sector A3.

Interlock techs tested safety systems in TeV sectors A-E, B0 crossover, Pbar and its transport lines.

Main Injector (MI) experts tested power supplies.

During the evening shift, operators interlocked TeV sectors A-E and the Transfer Hall, and the Pbar tunnels.

Operations established beam to Booster for studies, kicker scans, and injection tuning.

The Neutron Therapy Facility (NTF) took beam all shift.

Operators put the TeV separators into conditioning mode.

Saturday May 20

The midnight shift began with NTF conducting studies, and with the TeV separators conditioning.

Operators hipotted the TeV.

Interlock techs resumed safety system tests near the end of the midnight shift.

The scheduled power outage for Wilson hall began.

Power supply tests in the P1, P2, and P3 line uncovered two ground faults.

A Linac vacuum pump failed near the end of the day shift.

Sunday May 21

The midnight shift began with the TeV separator conditioning and Booster studies continuing.

NTF completed their studies.

Operators turned off the separators to hunt for a TeV ground fault. They found a possible cause with some cables that had been cut off and left touching a bus.

A Booster expert worked on the timing for the dump kickers and beam tuning.

Operators turned off MI to search for a ground fault and then resumed running.

Operations reported a power glitch around noon.

Monday May 22

The midnight shift began with TeV separator conditioning and MI ramping.

A Booster expert conducted studies.

The Plans for Today

The plans are to access MI to repair a beam valve. The TeV A-E and Transfer hall can be control accessed for work.

Machine Reports

Linac

No report.

Booster

The injection girder still needs work. The dump tune is off a little. The velocity meter needs repair. ORBUMP needs balancing and the Booster needs tuning to up the intensity.

MI

Techs need to inspect the beamstop. There might be a problem with water plating in P3. The MI signoff is complete. The P2 F1 house has a vacuum gauge problem that will need an access. Experts hope to get some beam this afternoon after the beam valve gets fixed.

Pbar

The Transport line is secure, but all other areas are in access. The AP1 line needs a vacuum pump down. The A30 bake was completed. Experts would like to test the Debuncher. They would like beam by Wednesday night.

TeV

All the separators are on and conditioning. The main bus was hipotted and faults found. Experts want a F sector access. Experts tested the Linux sequencers and found problems.

Recycler

Experts would like to start short batching protons. They also want to start electron studies. Sign offs should be completed today.

Switchyard 120

Experts are conducting walkthroughs today.

MiniBooNE

They report that they will be ready for beam.

NuMI

Experts are busy starting things up. There is still sump work going on. The chillers and fans are ready for leak tests. They will be ready for beam by the end of the week.

Collider Detector at Fermilab

Their LCW system is back on and they have begun flowing gas for their detector. They hope to have their detector closed up by Tuesday or Wednesday of next week.

D-Zero

They are checking their detector.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/22/2006    email Fermilab